1. Support Department Cost Allocation—Direct Method

Christmas Timber, Inc., produces Christmas trees. The trees are produced through a cutting and pruning process. Machine maintenance and janitorial labors are performed throughout the production process by non-production employees. Maintenance and janitorial costs are allocated based on machine hours used and the number of trees in each department, respectively. The company estimates that the cutting and pruning areas typically have about 6 and 54 trees, respectively, in them at 1 time. The company also estimates that the cutting process requires about 9 times as many machine hours as the pruning process. The total costs of each department are as follows:

Maintenance Department $6,000
Janitorial Department 4,000
Cutting Department 55,000
Pruning Department 13,000

Using the direct method of support department cost allocation, determine the total cost of each production department after allocating all support costs to the production departments.

Cutting
Department
Pruning
Department
Production departmentsʼ total costs $ $

2. Support department cost allocation—reciprocal services method

Davis Snowflake & Co. produces Christmas stockings in its Cutting and Sewing departments. The Maintenance and Security departments support the production of the stockings. Costs from the Maintenance Department are allocated based on machine hours, and costs from the Security Department are allocated based on asset value. Information about each department is provided in the following table:

Maintenance
Department
Security
Department
Cutting
Department
Sewing
Department
Machine hours 800    2,000 6,400    11,600
Asset value $2,000    $1,420 $2,500 $5,500
Department cost $42,480    $18,880 $62,000 $83,000

Determine the total cost of each production department after allocating all support department costs to the production departments using the reciprocal services method.

Cutting
Department
Sewing
Department
Production departmentsʼ total costs $ $
Answer #1

Answer to Question 1.

Allocation of Cost from Machining Department to Cutting and Pruning Department:

Ratio of allocation between Cutting and Pruning Department for Machining Cost = 9 : 1

Allocation of Machining Cost to Cutting Department = 9/ (9+1) * $6,000

Allocation of Machining Cost to Cutting Department = $5,400

Allocation of Machining Cost to Pruning Department = 1/ (9+ 1) * $6,000

Allocation of Machining Cost to Pruning Department = $600

Allocation of Cost from Janitorial Department to Cutting and Pruning Department:

Ratio of Allocation between Cutting and Pruning Department for Janitorial Cost = 6 : 54

Allocation of Janitorial cost to Cutting Department = 6 / (6 + 54) * $4,000

Allocation of Janitorial cost to Cutting Department = $400

Allocation of Janitorial cost to Pruning Department = 54/ (6 + 54) * $4,000

Allocation of Janitorial cost to Pruning Department = $3,600

Cutting Department Pruning Department
Production departments' total cost $5,800 $4,200